OpenStudy (crystal99):

A man lifts 20 bricks,each of weight 6N.What other information is needed to calculate the useful work done in lifting the bricks? A.the distance he lifts the bricks B.the mass of the bricks C.the time taken to lift the bricks D.the volume of the bricks

OpenStudy (anonymous):

The distance he lifts the bricks.

OpenStudy (crystal99):

Thank you for the answer.

OpenStudy (anonymous):

Your welcome. Happy to help :)

OpenStudy (anonymous):

how did u do this ..pls explain

OpenStudy (anonymous):

ok...thnx..silly me

OpenStudy (anonymous):

\[Work = Force \times Displacement \times \cos \theta\]

OpenStudy (crystal99):

Happy to know there are cool people here..:)